# Models That Prove Their Own Correctness

### v2

@@ A.1 Transcript Learning
The lower-bounding function is the agreement of the transcript generator induced by $P_{\theta}$ with the provided honest transcript generator $\mathcal{T}^{\ast}_{V}$. More formally, we let $\mathcal{T}^{\theta}_{V}$ denote the transcript generator induced by the model $P_{\theta}$: for each $x$, $\mathcal{T}^{\theta}_{V}(x)$ is simply the distribution over transcripts of interactions between $V$ and $P_{\theta}$ on input $x$. We first prove TL correctly estimates the gradient of $A(\theta)$ in its update step.
Lemma A.4 (TL gradient estimation).
@@
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. Fix an honest transcript generator $\mathcal{T}_{V}^{\ast}$. Let $\theta$ be the parameters of a model $P_{\theta}$ such that
@@
is differentiable in $\theta$. Then
$\nabla A(\theta)=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ \pi^{\ast}\sim\mathcal{T}^{\ast}_{V}\end{subarray}}\left[\alpha_{0}(\theta)\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta)\cdot\alpha_{s}(\theta)\right)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta)\right]$
where $S(r)$, $\beta_{r}(\theta)$, $\alpha_{s}(\theta)$ and $\vec{d}_{s}(\theta)$ are as defined in Algorithm 1.
Note that Lemma A.4 is true for any model $P_{\theta}$. Moreover, the random vector over which the expectation is taken (in the right hand side) is precisely the direction of the update performed in Algorithm 1. We now prove Lemma A.4, from which we derive Theorem 4.1.
Proof.
Throughout this proof, expectations and probabilities will be over the same distributions as in the lemma statement. First, by the law of total probability, and linearity of the gradient,
$\nabla A(\theta)\coloneqq\nabla_{\theta}\left(\Pr_{x,\pi^{*},\pi}\left[\pi=\pi^{*}\right]\right)=\nabla_{\theta}\left(\operatorname*{\mathbb{E}}_{x,\pi^{*}}\left[\Pr_{\pi}\left[\pi=\pi^{*}\right]\right]\right)=\operatorname*{\mathbb{E}}_{x,\pi^{*}}\left[\nabla_{\theta}\left(\Pr_{\pi}\left[\pi=\pi^{*}\right]\right)\right].$
Next, we use the law of total probability together with the autoregressive property of $P_{\theta}$ (Section 4) to switch from probabilities on transcripts, to products of next-token probabilities. Formally, consider any fixed input $x$, honest transcript $\pi^{\ast}=(y^{\ast},q_{1}^{\ast},a_{1}^{\ast},\dots,q_{R}^{\ast},a_{R}^{\ast})$, and denote a random transcript sampled from $\mathcal{T}_{V}^{\theta}(x)$ by $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})$. For any $r\in[R]$ denote the random variables $V_{q}^{<r}\coloneqq V_{q}(y,q_{1},a_{1},\dots,q_{r-1},a_{r-1})$ and $\mathcal{T}^{\theta,<r}_{V}\coloneqq\mathcal{T}^{\theta}_{V}(yq_{1}a_{1}\cdots a_{r-1}q_{r})$. Then,
$\Pr_{\pi}\left[\pi=\pi^{*}\right]$ $\coloneqq\Pr_{\pi}[(y^{\ast},q_{1}^{\ast},a_{1}^{\ast},\dots,q_{R}^{\ast},a_{R}^{\ast})=(y,q_{1},a_{1},\dots,q_{R},a_{R})]$ (3)
$=\Pr_{y\sim P_{\theta}(x)}[y=y^{\ast}]\cdot\prod_{\begin{subarray}{c}r\in[R]\end{subarray}}\Pr_{q\sim V_{q}^{<r}}[q=q_{r}^{\ast}]\cdot\Pr_{a\sim\mathcal{T}^{\theta,<r}_{V}}[a=a_{r}^{\ast}]$
$=\Pr_{y\sim P_{\theta}(x)}[y=y^{\ast}]\cdot\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\Pr_{q\sim V_{q}^{<r}}[q=q_{r}^{\ast}]\cdot\Pr_{\sigma\sim p_{\theta}(\pi^{\ast}_{<s})}[\sigma=\pi^{\ast}_{s}]$ (4)
$=\alpha_{0}(\theta)\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta)\cdot\alpha_{s}(\theta)\right),$ (5)
where Equation 3 uses independence of the verifier and model’s randomness, Equation 4 uses the autoregressive property of $P_{\theta}$ (Definition A.1), and Equation 5 is by definition of $\alpha_{s}$ and $\beta_{r}$.
Next, a basic calculus identity gives
Let us focus on the rightmost factor. By Equation 5,
$\nabla_{\theta}\log\left(\Pr_{\pi}\left[\pi=\pi^{*}\right]\right)$ $=\nabla_{\theta}\log\alpha_{0}(\theta)\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta)\cdot\alpha_{s}(\theta)\right)$
$=\nabla\log_{\theta}\alpha_{0}(\theta)+\sum_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\nabla_{\theta}\log\beta_{r}(\theta)+\nabla_{\theta}\log_{\theta}\alpha_{s}(\theta)$
$=\nabla\log_{\theta}\alpha_{0}(\theta)+\sum_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\nabla_{\theta}\log_{\theta}\alpha_{s}(\theta)$ (7)
$=\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\nabla_{\theta}\log_{\theta}\alpha_{s}(\theta)=\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta)$ (8)
where Equation 7 is because $\log\beta_{r}(\theta)\coloneqq\log\Pr_{q^{\prime}\sim V_{q}(x\pi^{\ast}_{<t})}[q^{\prime}=q]$ is a constant and therefore has a gradient of zeros, and Equation 8 is by definition of $\vec{d}_{s}(\theta)$. Combining Equations 5, 6 and 8 concludes the proof. ∎
Theorem A.5 (Theorem 4.1, formal).
Fix a verifier $V$, an input distribution $\mu$, and an autoregressive model family $\{P_{\theta}\}_{\theta\in\mathbb{R}^{d}}$, and a norm |||||$\cdot$||| | ⋅ | | on $\mathbb{R}^{d}$. Fix an honest transcript generator $\mathcal{T}^{\ast}_{V}$, and assume that the agreement function
is concave and differentiable in $\theta$. For any $\varepsilon>0$, let $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ be upper-bounds such that the following conditions hold.
@@
Denote by $\bar{\theta}$ the output of TL running for number of iterations
@@
Our strategy is to cast TL as Stochastic Gradient Ascent and apply B.2. Let $\varepsilon$, $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ as in the theorem statement be given. Let $\theta^{\ast}$ be such that $A(\theta^{\ast})\geq 1-\varepsilon/2$ and $||\theta^{\ast}||\leq B_{\mathrm{Norm}}$.
@@
$\left\lVert\alpha_{0}(\theta_{i})\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta_{i})\alpha_{s}(\theta_{i})\right)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i})\right\rVert\\ =\left|\alpha_{0}(\theta_{i})\cdot\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta_{i})\alpha_{s}(\theta_{i})\right|\cdot\left\lVert\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i})\right\rVert\\ \leq 1\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\left\lVert\vec{d}_{s}(\theta_{i})\right\rVert,$
where the inequality is because $\alpha_{s}(\theta_{i})$ and $\beta_{r}(\theta_{i})$ are probabilities, so $\leq 1$. Continuing, we have
@@
The first inequality is by definition of $B_{\mathrm{Lip}}$ as an upper-bound on the gradient of $P_{\theta}$’s logits. The second is because, by definition, $C$ is an upper bound on the number of tokens sent by the prover in the proof system, which is exactly the number of terms in the sum: $r$ indexes rounds, and $s$ indexes tokens sent in each round.
To conclude, Lemma A.4 shows that TL samples from a gradient estimator for $A(\theta)$, while the above equation shows that the gradient is upper-bounded by $C\cdot B_{\mathrm{Lip}}$. We can therefore apply B.2 to obtain
@@
A.2 Reinforcement Learning from Verifier Feedback
Our second learning method, Reinforcement Learning from Verifier Feedback (RLVF, Algorithm 2), does not require access to an honest transcript generator. Instead, the learner learns $P_{\theta}$ generates transcripts herself by emulating the interaction of the verifier with the current Self-Proving model $P_{\theta}$. When an accepting transcript is generated, the learner updates the parameters $\theta$ towards generating such transcript.
@@ A.2 Reinforcement Learning from Verifier Feedback
21:for $i=0,\dots,N-1$ do
2:Sample $x\sim\mu$.
3:Initialize $a_{0}\coloneqq y$ and $d_{i}\coloneqq\vec{0}$.
3      4:foreach Round of interaction $r=1,\dots R$ do
5:Sample the $r^{\text{th}}$ query # Emulate the verifier
$q_{r}\sim V_{q}(x,a_{0},q_{1},a_{1},\dots,q_{r},a_{r}).$
6:Sample the $r^{\text{th}}$ answer # Forwards pass
$a_{r}\sim P_{\theta}(x,a_{0},q_{1},a_{1},\dots,q_{r},a_{r},q_{a_{r+1}}).$
4            7:Let $\tau_{r}\coloneqq(a_{0},q_{1},\dots,a_{r-1},q_{r})$. for $s\in[L_{a}]$ do
5                  8:Let $a_{r,s}$ denote the $s^{\text{th}}$ token in $a_{r}$. Compute # Backwards pass
@@
6      9:if $V(x,y,q_{1},a_{1},\dots,q_{R},a_{R})\ \mathrm{accepts}$ then
7            10:Update
@@
11:Output $\bar{\theta}\coloneqq\frac{1}{N}\sum_{i\in[N]}\theta_{i}$.
@@
Secondly, in comparing Algorithms 1 and 2, we see that the latter (RLVF) does not keep track of the probabilities $\alpha_{s}$ and $\beta_{r}$. This is because, in RL terms, RLVF is an on-policy algorithm; it generates transcripts using the current learned model, unlike TL which samples them from a distribution whose parameterization is unknown to the learner. Hence, the update step in RLVF is simpler than TL. Furthermore, the RLVF learner does not require access to the density function of the query generator $V_{q}$ (Definition A.2) unlike its TL counterpart.
We now prove that the update step in RLVF maximizes the Verifiability of $P_{\theta}$; this is analogous to Lemma A.4 for TL. We leave it for future work to use Lemma A.6 to obtain convergence bounds on RLVF (analogous to Theorem A.5). As mentioned in Section 4.2, the gap between the lemma and a full convergence theorem (informally) reduces to the problem of obtaining convergence bounds for Policy Gradient methods, a challenging and active research direction (e.g. Agarwal et al. 2021).
Lemma A.6 (RLVF gradient estimation).
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. For any transcript $(x,y,q_{1},\dots,a_{R})$ we let $\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})$ denote the indicator random variable which equals 1 if and only if $V$ accepts the transcript. For any model $P_{\theta}$, denote by $\mathrm{ver}(\theta)$ the verifiability of $P_{\theta}$ with respect to $V$ and $\mu$ (Definition 3.4). For any $\theta$, if $\mathrm{ver}(\theta)$ is differentiable in $\theta$, then
$\nabla\mathrm{ver}(\theta)=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\\ (q_{r},a_{r})_{r=1}^{R}\end{subarray}}\left[\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in[L_{a}]\end{subarray}}\vec{d}_{s}(\theta)\right]$
@@
Recall the transcript generator of $P^{\theta}$, denoted by $T^{\theta}_{V}$ (see Lemma A.4). By the definitions of Verifiability in Definition 3.4 and $V(x,y,q_{1},\dots,a_{R})$ in the lemma statement,
@@
$=\operatorname*{\mathbb{E}}_{x\sim\mu}\left[\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}}[\mathrm{Acc}_{V}(x,\pi)\ ]\right]$ (9)
@@
$\Pi^{\ast}(x)\coloneqq\left\{\pi^{\ast}\in\Sigma^{\ast}:\mathrm{Acc}_{V}{x,\pi^{\ast}}\ \mathrm{accepts}\right\}.$
Noting that $\Pi^{\ast}(x)$ has finite or countably infinite cardinality, for any fixed input $x$ we can write
$\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}}[\mathrm{Acc}_{V}(x,\pi)]=\sum_{\pi^{\ast}\in\Pi^{\ast}(x)}\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}[\pi=\pi^{\ast}].$ (10)
We will use Equations (3) through (8) in the proof of Lemma A.4. Up to a change in index notation, these show that, for any $\pi^{\ast}$,
@@
Combining Equations 9 and 10, by linearity of expectation we have that
$\nabla_{\theta}\mathrm{ver}(\theta)$ $=\sum_{\pi^{\ast}\in\Pi^{\ast}(x)}\nabla_{\theta}\Pr_{\pi\sim\mathcal{T}^{\theta}(x)}[\pi=\pi^{\ast}]$
@@
Appendix B Preliminaries on Stochastic Gradient Ascent

### v3

@@ 9      Update
Figure 4: Transcript Learning, visualized. To understand Algorithm 1, consider the above visualization. In Phase 1, $N$ honest transcripts are collected by letting an Honest Prover interact with the Verification Algorithm; these will be the samples from the honest transcript generator $\mathcal{T}^{\ast}_{V}(x)$. Phase 2 describes the execution of Algorithm 1 itself: For each honest transcript $\pi^{\ast}$ (lines 2-3), and for each prefix $\pi_{s}$ of this transcript (lines 4-6), the $\alpha_{s}(\theta_{i})$ and $\vec{d_{s}}(\theta_{i})$ are computed via forwards and backwards passes, respectively (line 7). After iterating through all prefixes, the parameters $\theta_{i}$ are updated (line 8).
@@
The lower-bounding function is the agreement of the answers generated by $P_{\theta}$ with the answers provided by the honest transcript generator $\mathcal{T}^{\ast}_{V}$. More formally, we let $\mathcal{T}^{\theta}_{V}$ denote the transcript generator induced by the model $P_{\theta}$ when interacting with $V$: for each $x$, $\mathcal{T}^{\theta}_{V}(x)$ is the distribution over transcripts of interactions between $V$ and $P_{\theta}$ on input $x$. We stress that $\pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)$ and $\pi\sim\mathcal{T}^{\theta}_{V}(x)$ are transcripts produced when interacting with the same verifier queries; we can think of the verifier as simultaneously interacting with the honest prover and with the model $P_{\theta}$. [^10] The way it is presented in the algorithm (and implemented in the experiments), first the verifier is called by $\mathcal{T}^{\ast}_{V}$ and outputs queries $(q^{*}_{1},\dots q^{*}_{R})$, and then the model is prompted with the verifier queries one a time. This maintains soundness, since a proof system is sound as long as the prover does not know the verifier’s queries in advance. In what follows, we use $\pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)$ and $\pi\sim\mathcal{T}^{\theta}_{V}(x)$ to denote two transcripts that share the same queries. That is, taking $\pi^{\ast}=(y^{\ast},q_{1}^{\ast},a_{1}^{\ast},\dots,q_{R}^{\ast},a_{R}^{\ast})$ to denote an accepting transcript sampled from $\mathcal{T}^{\ast}_{V}(x)$, and $\pi=(y,q^{*}_{1},a_{1},\dots,q^{*}_{R},a_{R})$ to denote a random transcript sampled from $\mathcal{T}_{V}^{\theta}(x)$, we say that $\pi$ and $\pi^{*}$ agree if they agree on the prover answers, namely if:
$(y,a_{1},\dots,a_{R})=(y^{\ast},a_{1}^{\ast},\dots,a_{R}^{\ast}).$
This definition implicitly uses the independence of the verifier and model’s randomness. We first prove that TL correctly estimates the gradient of $A(\theta)$ in its update step.
Lemma B.4 (TL gradient estimation).
@@
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. Fix an honest transcript generator $\mathcal{T}_{V}^{\ast}$. Let $\theta$ be the parameters of a model $P_{\theta}$ and let
$A(\theta)\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ \pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)\\ \pi\sim\mathcal{T}^{\theta}_{V}(x)\end{subarray}}\left[\pi=\pi^{*}\right].$
Then,
$\nabla A(\theta)=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ \pi^{\ast}\sim\mathcal{T}^{\ast}_{V}\end{subarray}}\left[\prod_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\alpha_{s}(\theta)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta)\right]$
where $S(r)$, $\alpha_{s}(\theta)$ and $\vec{d}_{s}(\theta)$ are as defined in Algorithm 1.
Note that Lemma B.4 is true for any model $P_{\theta}$. Moreover, the random vector over which the expectation is taken (in the right hand side) is precisely the direction of the update performed in Algorithm 1. We now prove Lemma B.4, from which we derive Theorem 4.1.
Proof.
Throughout this proof, expectations and probabilities will be over the same distributions as in the lemma statement. First, we use the law of total probability together with the autoregressive property of $P_{\theta}$ (Section 4) to switch from probabilities on transcripts, to products of next-token probabilities. Formally, consider a fixed input $x$, an honest transcript $\pi^{\ast}=(y^{\ast},q_{1}^{\ast},a_{1}^{\ast},\dots,q_{R}^{\ast},a_{R}^{\ast})$, and denote a random transcript sampled from $\mathcal{T}_{V}^{\theta}(x)$ when using the same verifier queries by $\pi=(y,q_{1}^{\ast},a_{1},\dots,q_{R}^{\ast},a_{R})$. For any $r\in[R]$ denote the random variable $\mathcal{T}^{\theta,<r}_{V}\coloneqq\mathcal{T}^{\theta}_{V}(yq^{\ast}_{1}a_{1}\cdots a_{r-1}q^{\ast}_{r})$. Then,
$\Pr_{\pi}\left[\pi=\pi^{*}\right]$ $=\Pr_{\pi}[(y,a_{1},\dots,a_{R})=(y^{\ast},a_{1}^{\ast},\dots,a_{R}^{\ast})]$ (3)
$=\Pr_{y\sim P_{\theta}(x)}[y=y^{\ast}]\cdot\prod_{\begin{subarray}{c}r\in[R]\end{subarray}}\Pr_{a\sim\mathcal{T}^{\theta,<r}_{V}}[a=a_{r}^{\ast}]$
$=\Pr_{y\sim P_{\theta}(x)}[y=y^{\ast}]\cdot\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\Pr_{\sigma\sim p_{\theta}(\pi^{\ast}_{<s})}[\sigma=\pi^{\ast}_{s}]$ (4)
$=\prod_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\alpha_{s}(\theta),$ (5)
where, as noted above, Equation 3 uses the independence of the verifier and model’s randomness, Equation 4 uses the autoregressive property of $P_{\theta}$ (Definition B.1), and Equation 5 is by definition of $\alpha_{s}$ and of $a_{0}$. Next, a basic calculus identity gives
This implicitly assumes that $\Pr_{\pi}\left[\pi=\pi^{*}\right]$ is differentiable in $\theta$; indeed, this follows from Definition B.1, where the logits of the model were assumed to by differentiable. Let us focus on the rightmost factor. By Equation 5,
$\nabla_{\theta}\log\left(\Pr_{\pi}\left[\pi=\pi^{*}\right]\right)$ $=\nabla_{\theta}\log\left(\prod_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\alpha_{s}(\theta)\right)=\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\nabla_{\theta}\log\alpha_{s}(\theta)=\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta)$ (7)
where the last equality is by definition of $\vec{d}_{s}(\theta)$. Combining Equation 5 and Equation 6 gives
$\nabla_{\theta}\left(\Pr_{\pi}\left[\pi=\pi^{*}\right]\right)=\prod_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\alpha_{s}(\theta)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta).$
By the law of total probability and the linearity of the gradient,
$\operatorname*{\mathbb{E}}_{x,\pi^{*}}\left[\nabla_{\theta}\left(\Pr_{\pi}\left[\pi=\pi^{*}\right]\right)\right]=\nabla_{\theta}\left(\operatorname*{\mathbb{E}}_{x,\pi^{*}}\left[\Pr_{\pi}\left[\pi=\pi^{*}\right]\right]\right)=\nabla_{\theta}\left(\Pr_{x,\pi^{*},\pi}\left[\pi=\pi^{*}\right]\right)=\nabla_{\theta}A(\theta).$
which concludes the proof. ∎
Theorem B.5 (Theorem 4.1, formal).
Fix a verifier $V$, an input distribution $\mu$, an autoregressive model family $\{P_{\theta}\}_{\theta\in\mathbb{R}^{d}}$, and a norm |||||$\cdot$||| | ⋅ | | on $\mathbb{R}^{d}$. Fix an honest transcript generator $\mathcal{T}^{\ast}_{V}$, and assume that the agreement function
@@
is concave in $\theta$, where the verifier queries are the same in $\pi^{*}$ and $\pi$. For any $\varepsilon>0$, let $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ be upper-bounds such that the following conditions hold.
@@
Denote by $\bar{\theta}$ the output of TL running for number of iterations $N$ where
@@
Our strategy is to cast TL as Stochastic Gradient Ascent and apply C.2. Let $\varepsilon$, $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ as in the theorem statement be given. Let $\theta^{\ast}$ be such that $A(\theta^{\ast})\geq 1-\varepsilon/2$ and $||\theta^{\ast}||\leq B_{\mathrm{Norm}}$.
@@
$\left\lVert\prod_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\alpha_{s}(\theta_{i})\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i})\right\rVert=\left|\prod_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\alpha_{s}(\theta_{i})\right|\cdot\left\lVert\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i})\right\rVert\\ \leq 1\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\left\lVert\vec{d}_{s}(\theta_{i})\right\rVert,$
where the inequality is because $\alpha_{s}(\theta_{i})$ are probabilities, so $\leq 1$. Continuing, we have
@@
The first inequality is by definition of $B_{\mathrm{Lip}}$ as an upper-bound on the gradient of $P_{\theta}$’s logits. The second is because, by definition, $C$ is an upper-bound on the number of tokens sent by the prover in the proof system, which is exactly the number of terms in the sum: $r$ indexes rounds, and $s$ indexes tokens sent in each round.
To conclude, Lemma B.4 shows that TL samples from a gradient estimator for $A(\theta)$, while the above equation shows that the gradient is upper-bounded by $C\cdot B_{\mathrm{Lip}}$. We can therefore apply C.2 to obtain
@@
B.2 Reinforcement Learning from Verifier Feedback
Our second learning method, Reinforcement Learning from Verifier Feedback (RLVF, Algorithm 2), does not require access to an honest transcript generator. Instead, the learner generates transcripts herself by emulating the interaction of the verifier with the current Self-Proving model $P_{\theta}$. When an accepting transcript is generated, the learner updates the parameters $\theta$ towards generating such transcript.
@@ B.2 Reinforcement Learning from Verifier Feedback
2for $i=0,\dots,N-1$ do
3       Sample $x\sim\mu$.
4       Initialize $a_{0}\coloneqq y\sim P_{\theta_{i}}(x)$.
5       foreach Round of interaction $r=1,\dots R$ do
6             Sample the $r^{\text{th}}$ query # Emulate the verifier
$q_{r}\sim V_{q}(x,a_{0},q_{1},a_{1},\dots,q_{r-1},a_{r-1}).$
Sample the $r^{\text{th}}$ answer # Forwards pass
$a_{r}\sim P_{\theta_{i}}(x,a_{0},q_{1},a_{1},\dots,q_{r}).$
Let $\tau_{r}\coloneqq(a_{0},q_{1},\dots,a_{r-1},q_{r})$.
7             for $s\in[L_{a}]$ do
8                   Let $a_{r,s}$ denote the $s^{\text{th}}$ token in $a_{r}$. Compute # Backwards pass
@@ 6             Sample the $r^{\text{th}}$ query # Emulate the verifier
9      if $V(x,y,q_{1},a_{1},\dots,q_{R},a_{R})\ \mathrm{accepts}$ then
10             Update
@@
Output $\bar{\theta}\coloneqq\frac{1}{N}\sum_{i\in[N]}\theta_{i}$.
@@
Secondly, in comparing Algorithms 1 and 2, we see that the latter (RLVF) does not keep track of the probabilities $\alpha_{s}$. This is because, in RL terms, RLVF is an on-policy algorithm; it generates transcripts using the current learned model, unlike TL that samples them from a distribution whose parameterization is unknown to the learner. Hence, the update step in RLVF is simpler than TL.
We now prove that the update step in RLVF maximizes the Verifiability of $P_{\theta}$; this is analogous to Lemma B.4 for TL. We leave it for future work to use Lemma B.6 to obtain convergence bounds on RLVF (analogous to Theorem B.5). As mentioned in Section 4.2, the gap between the lemma and a full convergence theorem (informally) reduces to the problem of obtaining convergence bounds for Policy Gradient methods, a challenging and active research direction (e.g. Agarwal et al. 2021).
Indeed, the update step that the algorithm takes can be expressed as the random vector over which the expectation is taken (in the right hand side).
Lemma B.6 (RLVF gradient estimation).
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. For any transcript $(x,y,q_{1},\dots,a_{R})$ we let $\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})$ denote the indicator random variable which equals 1 if and only if $V$ accepts the transcript. For any model $P_{\theta}$, denote by $\mathrm{ver}(\theta)$ the verifiability of $P_{\theta}$ with respect to $V$ and $\mu$ (Definition 3.4). Then, for any $\theta$,
$\nabla_{\theta}\mathrm{ver}(\theta)=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\\ (q_{r},a_{r})_{r=1}^{R}\end{subarray}}\left[\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in[L_{a}]\end{subarray}}\vec{d}_{s}(\theta)\right]$
@@
Recall the transcript generator of $P_{\theta}$, denoted by $\mathcal{T}^{\theta}_{V}$ (see Lemma B.4). By the definitions of Verifiability in Definition 3.4 and $V(x,y,q_{1},\dots,a_{R})$ in the lemma statement,
@@
$=\operatorname*{\mathbb{E}}_{x\sim\mu}\left[\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}[\mathrm{Acc}_{V}(x,\pi)\ ]\right]$ (8)
@@
$\Pi^{\ast}(x)\coloneqq\left\{\pi^{\ast}\in\Sigma^{\ast}:\mathrm{Acc}_{V}({x,\pi^{\ast}})=1\right\}.$
We can assume that $\Pi^{\ast}(x)$ has finite cardinality, since $V$’s running time is bounded and hence the number of different transcripts that it can read (and accept) is finite. For any fixed input $x$, we can express its acceptance probability by the finite sum:
$\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}[\mathrm{Acc}_{V}(x,\pi)]=\sum_{\pi^{\ast}\in\Pi^{\ast}(x)}\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}[\pi=\pi^{\ast}].$ (9)
We will use Equations (3) through (7) in the proof of Lemma B.4. Up to a change in index notation, these show that, for any $\pi^{\ast}$,
@@
Combining Equations 8 and 9, by linearity of expectation we have that
$\nabla_{\theta}\mathrm{ver}(\theta)$ $=\operatorname*{\mathbb{E}}_{x\sim\mu}\left[\sum_{\pi^{\ast}\in\Pi^{\ast}(x)}\nabla_{\theta}\Pr_{\pi\sim\mathcal{T}^{\theta}(x)}[\pi=\pi^{\ast}]\right]$
@@
Appendix C Preliminaries on Stochastic Gradient Ascent